OpenStudy (anonymous):

A 30 ft. ladder reaches a height of 24 ft. when it is placed against a building and a 25 ft. ladder reaches 20 ft. The base of which ladder is further from the base of the building and by how many feet?

OpenStudy (campbell_st):

well you could assume they form right angled triangles... and use pythagoras the ladders are the hypotenuses... and you would be finding a shorter side in both triangles.

OpenStudy (anonymous):

thats what we did in class but im really confused

OpenStudy (campbell_st):

ok...  use the triangles I've drawn

OpenStudy (anonymous):

so the actual height of the ladder doesn't matter

OpenStudy (anonymous):

@campbell_st

OpenStudy (campbell_st):

well the height of the ladder is the hypotenuse of the right triangles so using pythagoras you have \[x^2 = 30^2 - 25^2\] and \[y^2 = 24^2 - 20^2\]

OpenStudy (campbell_st):

when you get the answers, take the square root of both to find out how far the base of each ladder is from the wall..

OpenStudy (anonymous):

ok i got it now

OpenStudy (anonymous):

thx